If 237 g of H2 is mixed with N2 at a pressure of 152,000 torr and a temperature of 400°C, what volume of N2 gas is needed for the reaction?
1
Expert's answer
2020-04-13T02:34:32-0400

n (H2) = m/M = 237 g / 2.02 g/mol = 117.33 mol

pV = nRT

V (H2) = nRT / P

T = 400 + 273 = 673 K

P = 152,000 torr = 20,265,000 Pa

V (H2) = 117.33 mol * 8.314 J/mol*K * 673 K / 20,265,000 Pa

V (H2) = 0.03 m3

V (N2) = 1/3 * 0.03 m3 = 0.01 m3 according to the equation:

N2 + 3H2 ⇄ 2NH3